A rectangular loop of wire that can rotate about an axis through its center is placed between the poles of a magnet in a magnetic field with a strength of 0.357 T (with the magnetic field parallel to the plane of the loop). The length of the loop L is 15.5 cm and its width w is 10.0 cm. What emf is developed in the loop if the loop is rotated form this position, in 1.00 ms, to a position where the plane of the loop makes an angle of 54.5 degrees with the magnetic field.
1
Expert's answer
2019-03-08T07:13:14-0500

Question:


A rectangular loop of wire that can rotate about an axis through its center is placed between the poles of a magnet in a magnetic field with a strength of 0.357 T (with the magnetic field parallel to the plane of the loop). The length of the loop L is 15.5 cm and its width w is 10.0 cm. What emf is developed in the loop if the loop is rotated form this position, in 1.00 ms, to a position where the plane of the loop makes an angle of 54.5 degrees with the magnetic field.


Solution:

According to the Faraday law emf E = dФ/dt ≈ ΔФ/Δt = LwBcos(90-54.5)/Δt = 1.55*3.57*cos35.5 = 4.5 (V).


The answer:


The emf E = 4.5 V.